OpenStudy (anguyennn):

(x-1)^2-13=3, how do you find the x?

OpenStudy (anonymous):

add 13 take the square root, don't forget the \(\pm\) then add 1

OpenStudy (anguyennn):

what?

OpenStudy (anonymous):

add 13 to both sides what do you get?

OpenStudy (anguyennn):

16

OpenStudy (anonymous):

actually you get \[(x-1)^2=16\]

OpenStudy (anguyennn):

yes

OpenStudy (anonymous):

can you think of a number whose square is 16?

OpenStudy (anguyennn):

4

OpenStudy (anonymous):

ok so we can solve \(x-1=4\) easily enough now can you think of another number whose square is 16 (hint, it is negative)

OpenStudy (anguyennn):

-4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you have two equations to solve \[x-1=4\] and also \[x-1=-4\]

OpenStudy (anguyennn):

so will x be x=-5 and x=3

OpenStudy (anonymous):

no

OpenStudy (anonymous):

to solve \[x-1=4\]you need to ADD 1 to both sides

OpenStudy (anguyennn):

5

OpenStudy (anguyennn):

and -3